OpenStudy (clongoria2):

Medal and Fan for answer! When you are doing exponential functions, how would you go about solving this? 4^x+3. How do you involve the three in this equation. Say we plugged in -3 as the x. 4^-3+3. What would be the answer?

OpenStudy (michele_laino):

please note that: \[4^{-3}+3=\frac{ 1 }{ 4^{3} }+3=\frac{ 1 }{ 64 }+3=\frac{ 1+64*3 }{ 64 }=\] \[=\frac{ 1+192 }{ 64 }=\frac{ 193 }{ 64 }\]

OpenStudy (clongoria2):

How does it jump from that?

OpenStudy (michele_laino):

when you have a power with a negative exponent, then that power goes to denominator with the same exponent but reversed sign. for example: \[2^{-5}=\frac{ 1 }{ 2^{5} }\]

OpenStudy (clongoria2):

How do you convert a fraction back into an exponent?

OpenStudy (clongoria2):

Like how you have 193/64. Is it possible to convert this back into an exponent?

OpenStudy (michele_laino):

in general no, it is necessary to perform the factorization in prime numbers. for example, I know that: \[64=2^{6}\] and 193 is a prime number, so without any other infromation, your answer is negative

OpenStudy (clongoria2):

Thanks, that helped alot.

OpenStudy (michele_laino):

Thanks!